OpenStudy (anonymous):

how to find csc(x)=sec(x) for x from 0,2pi

OpenStudy (anonymous):

\[\frac{1}{sinx} = \frac{1}{cosx}\] multiply by sinx remember: sin/cos = tan tanx = 1 solve this in the given region and you have your values of x

OpenStudy (anonymous):

im happy to check your answers so feel free to post them if you want

OpenStudy (anonymous):

yeah why cant i do sinx=cosx sinx=sin(pi/2-x) .:x=pi/2-x 2x=pi/2 x=pi/4 missing soln?

OpenStudy (anonymous):

yeah missing solutions, i wouldn't reccomend that method, although it is valid. this is the correction: you're using cos(x) = sin( (pi/2) -x) but we could just as well use cos(x) = sin( (5pi/2) -x) and get the other solution

OpenStudy (anonymous):

recommend*